OpenStudy (crashonce):

3/4cos20−3√2sin20 = 1/8 prove this

OpenStudy (crashonce):

\[\frac{ 3 }{4}\cos20 - \frac{ \sqrt{3} }{ 2 }\sin 20 = \frac{ 1 }{8}\]